Find the center of mass of each of the following: (a) Athin wire bent into the form of a three-sided, block-shaped "Ll"with each segment of equal length b (b) A quadrant of a uniform circular lamina of radius b (c) The area bounded by parabola y =x2/b and the line y =b (d) The volume bounded by paraboloid of revolution z = (x2 +y2)/b and the plane z = b 


1
Expert's answer
2021-12-15T16:21:57-0500

a)

"x=\frac b2,~y=\frac b3."

b)

"x=0,~y=0."

c)

"x=0,~y=\frac {b\sqrt 2}2."

d)

"x=0,~y=0,~z=\frac{b\sqrt 3}2."